If car A is at 40 km/h and car B is at 10 km/h in the opposite direction, what is the velocity of the car A relative to the car B


1
Expert's answer
2021-05-19T17:03:11-0400

The velocity of the car A relative to the car B is


"v_{AB}=v_A+v_B=40+10=50\text{ km/h}"

if A and B mowe away from each other.